OpenStudy (anonymous):

In our new car, we were able to drive 30 miles in a half an hour. From this information, we can determine the car's

OpenStudy (austinl):

Rate?

OpenStudy (austinl):

@texaschic101 \(d=rt\) Where, \(d=\text{distance}\Rightarrow30mi\\ r=\text{rate}\\ t=\text{time}\Rightarrow30min\) \(30mi=r(30min)\) \(\displaystyle r=\frac{d}{t}=\frac{30miles}{0.5hours}=60mph\)

OpenStudy (texaschic101):

I think you are correct austinL....that makes more sense then my answer so I deleted mine. You got this one :)